# First Beam Neutrinos Observed with an LAPPD in the ANNIE Experiment

### v1

@@ Abstract
The Accelerator Neutrino Neutron Interaction Experiment (ANNIE) probes the physics of neutrino-nucleus interactions in a gadolinium-loaded water (Gd-water) target while serving as a flexible testbed for advanced next-generation optical neutrino detection technologies. These advanced technologies include novel detection media (particularly Gd-water and hybrid Cherenkov-scintillation through water-based liquid scintillator) and novel photosensors. In this paper we demonstrate the first implementation of a fully-integrated setup for Large Area Picosecond PhotoDetectors (LAPPDs) in a neutrino experiment. Details are presented regarding the design, commissioning, and deployment of an LAPPD and the supporting systems. We also present the first neutrino interactions ever observed with an LAPPD.
@@ 1 Introduction
LAPPDs are MicroChannel Plate (MCP)-based [1] imaging photosensors with a variety of potential applications in high-energy and nuclear physics, as well as practical applications in medical imaging [2] and other areas. Produced using advanced fabrication techniques such as atomic layer deposition, LAPPDs offer gain [3], approximately 50 ps time resolution [4], and unprecedented large area coverage. At 20 cm x 20 cm, they have a form factor nine times larger than the next-largest MCP-PMTs, at significantly lower cost per unit area. The large LAPPD area is coupled with a unique ability to correlate photon arrival times with sub-cm position resolution, determined primarily by the structure of the anode.
Until this point, studies of LAPPDs have been limited to proof-of-principle studies on the benchtop, focused primarily on characterizing LAPPD response in the single-photoelectron regime. Investigating LAPPD performance in the context of a running experiment, using actual neutrino interactions, represents a critical milestone toward establishing LAPPDs as viable next-generation photodetector for particle physics applications. Here we present the first neutrino interactions detected using an LAPPD in a working experiment and document the relevant technical details of the system developed to further the application readiness of this technology. In Section 2, we briefly describe the ANNIE detector and the experimental context. In Sec. 3, we describe the integration of an LAPPD into the ANNIE experiment. Finally, in Section 4, we establish the first detection of beam neutrinos with an LAPPD.
@@ 2 Review of the ANNIE Experiment
ANNIE is both a physics experiment studying neutrino-nucleus interactions in a water target and a unique US platform for developing next-generation neutrino detection technology [5]. The primary physics goal of ANNIE is to study neutrino interactions on water, with specific focus on measuring the abundance of final state neutrons from neutrino-nucleus interactions, using gadolinium (Gd) enhanced water. Measurements of final-state neutron multiplicity will improve our understanding of neutrino-nucleus interactions. The complex many-body physics of neutrino-nucleus scattering is among the dominant systematic uncertainties in precision long-baseline oscillation measurements [6, 7].
ANNIE also supports a broad R&D program that serves the experiment’s core physics goals while developing and testing new technologies for the future. One of the central goals of ANNIE is the first demonstration of LAPPDs in a neutrino experiment, as described in this paper. Other milestones of the R&D program include deployment of the first Gd-loaded water target in a neutrino beam (2021) and the first water-based liquid Scintillator (WbLS) target in a neutrino beam [8].
@@ 2.1 The ANNIE Detector
Figure 1 shows the ANNIE detector configuration. The main target consists of an upright cylindrical steel tank filled with 26 tons of Gd-loaded (0.1% by mass) ultra-pure deionized water. The photodetectors inside the tank are a mix of 8-, 10- and 11-inch conventional photomultiplier tubes (PMTs) that provide coverage in all directions. Eight regularly spaced access ports around the lid of the water tank each permit a column of up to three LAPPDs to be inserted into the tank. Data presented in this paper was acquired with a single LAPPD at a central position on wall of the tank downstream of the beam. In addition to the water volume, the full ANNIE detector system also includes a downstream muon detection system (refurbished from SciBooNE’s Muon Range Detector, or MRD [9]) that provides the primary constraint on muon energy and direction, as well as a scintillator-based Front Muon Veto (FMV) upstream of the water target.
A muon produced by a neutrino interaction in the fiducial volume will propagate through the tank, producing Cherenkov light along its path, before entering into and passing through the layers of scintillator paddles and iron absorber in the MRD. The position and timing of scintillator hits, combined with information about the Cherenkov light distribution from the tank photodetectors, are used to reconstruct the muon track. Neutrons produced by the neutrino interaction scatter and lose energy (thermalize) allowing them to be captured on either H or Gd in the active volume. The Gd loading substantially increases the neutron detection efficiency by increasing the capture cross-section for thermal neutrons (thereby shortening the neutron capture time) and providing a higher-energy, more easily-detected 8 GeV de-excitation -ray cascade [10, 11].
@@ 2.2 Booster Neutrino Beam
The ANNIE detector and data acquisition system are designed around the Booster Neutrino Beam (BNB) [12], which delivers neutrinos in discrete bursts timed to the accelerator Radio Frequency (RF) structure. The low duty cycle of the BNB allows the experiment to statistically suppress continuum backgrounds by confining the readout of the various detector sub-systems to a short time window around the arrival of neutrinos. A subset of off-beam data provides a sample of the continuum backgrounds, while laser, LED, and radioactive sources are used for calibration.
BNB neutrinos are produced from the decay of mesons made by colliding 8 GeV protons with a stationary beryllium target located at the MI-12 facility. Pions in a narrow range of momenta are selected using magnetic horns and steered to a decay pipe, where they decay to form a beam of 93 pure muon neutrinos (with 6 muon anti-neutrino and 1 electron (anti) neutrinos) with an energy spectrum peaked at a little below 600 MeV [13].
@@ 2.3 System Timing and DAQ Software
The system clock for the entire ANNIE experiment is provided by a CAEN V1495 general purpose VME board with a customizable FPGA, dubbed the ANNIE Central Trigger Card (CTC) [14]. The CTC is disciplined to a 25 MHz clock signal provided by a customized Spectrum Instruments TM-4 GPS clock with a double-ovenized crystal oscillator [15]. The CTC multiplies up the 25 MHz input clock to 125 MHz and passes that clock to the various subsystems. All subsystems match their timing to this 125 MHz clock via phase-locked loops (PLLs). Where needed, the CTC forwards a heartbeat, or pulse-per-second (PPS) signal, also provided by the GPS clock, to the various subsystems.
The CTC also manages, forwards, and timestamps a variety of beam-related and calibration triggers. Two of these signals are of particular importance. The CTC forms a beam trigger or beam gate as the logical and of a fiberoptic copy of the Booster Extraction Sync (BES) signal which precedes beam extraction by 320 s, and the Tevatron clock $1D signal, which indicates that the beam will be accelerated and extracted to the BNB and hence delivered to ANNIE. It then forwards time-delayed copies of this beam trigger to the readout electronics for the conventional PMT readout system and MRD, triggering both systems to read out all data related to a spill.
@@
For DAQ software, ANNIE uses the ToolDAQ framework, developed at Warwick University for the HyperK collaboration [16]. Asynchronous modular code sequences (ToolChains) pull data from each of the subsystems. An overall system loop on the top-level DAQ computer queries and pulls data from each subsystem ToolChain when available. CTC data, which includes timestamps for the beam input and heartbeat signals, is also pulled as a data stream. These time stamps are used to tie together the responses from individual subsystems to build a complete record of the detector’s response to a beam spill, cosmic ray, or calibration source, hereafter referred to as an event.
@@ 3 Technical Integration of an LAPPD into ANNIE
A number of technical and practical challenges were addressed in moving from benchtop studies of LAPPDs to their active use in a running experiment. Achieving time resolutions better than 100 ps required both fast sampling electronics and synchronization of the LAPPDs and the other ANNIE subsystems with a reliable central clock. The power dissipation from the readout electronics had to be minimized to reduce heating effects at the LAPPD and in the interior of the waterproof housing. This ensured stable operating conditions (the resistance and performance of the MCPs can be temperature-sensitive) and also minimized temperature-related physical stresses on the LAPPD. Signals and power were routed between the surface and the tank in a way that maximized signal fidelity and power stability while minimizing the risk of flooding the housing.
@@ 3.1 Surface Electronics
VME-based Annie Central Cards (ACCs) provide communications and clock synchronization with the underwater readout system, and serve as conduits for transferring the digital outputs of the LAPPDs to the ANNIE DAQ computer using a Raspberry Pi. Each ACC card is capable of servicing up to 2 LAPPDs. The ACCs are controlled by Altera Arria V FPGAs [17]. The communication link connecting the ACC and ACDC boards relies on the Low Voltage Differential Signaling (LVDS) logic standard. Each link between an ACC and ACDC consists of 4 twisted pairs: two lines for transmitting and receiving data, one line for sending the beam gate, and a clock signal for ensuring clock-domain alignment. A USB-serial buffer chip operating at 48 Mbps transfers data between the ACC and the control computer. While the ACC cards are designed to sit in a VME crate and have a standard VME form-factor, they do not require a VME backplane. They are housed in a VME crate with no backplane and powered by an Excelsys power supply[18].
@@ 3.2 Packaged ANNIE LAPPD (PAL)
A PAL consists of a waterproof housing containing the LAPPD assembly with readout electronics (ACDC cards), the slow controls board (LVHV), and independent leak-sensing and temperature-monitoring devices, as well as floodproof connectors that mate to waterproof cables. These elements are pictured in Figure 3.
@@
The first LAPPD deployed in ANNIE, LAPPD-40 (numbered reflecting its production order by the manufacturer, Incom Inc.), is a first-generation model LAPPD [3] instrumented with 28 silk-screened silver anode microstrips that are integral with the module. Photon hits are reconstructed based on the arrival of signal pulses at the two ends of each microstrip, as detailed further in Ref [19]. The nominal gain of LAPPD-40, as measured by Incom, is at 950V per MCP, and the average Qauntum Efficiency (QE) is 17. This performance is on the low end of the LAPPDs received by ANNIE, due in part to LAPPD-40 being one of the earlier functioning models.
@@ 3.2.1 Structure of the PAL
The physical waterproof housing of the PAL consists of a 2-inch thick PVC sidewall sandwiched between a stainless steel backplate and a quarter-inch thick UVT acrylic window. The window and backplate are bolted onto the sidewall with chemical-resistant Viton fluoroelastomer 3/16-inch O-rings [20] providing the waterproof seal. All of the power and communications lines connect to the LVHV board, which is affixed to the inside top of the sidewall. The LAPPD assembly is mounted onto tabs extending from the sidewall. All HV cables, power and communications lines from the ACDC cards and trigger cards plug into the LVHV board.
@@ 3.2.2 Digitization and Triggering in the PAL
Sampling and digitization of LAPPD waveforms is handled by ACDC cards, custom boards designed at the University of Chicago, that attach to the APB through high-density connectors. The ACDCs are built around the 6-channel PSEC4 waveform sampling ASIC, which provides 10 GS/s sampling rates and 1 GHz analog bandwidth [21]. Each ACDC card contains five PSEC4 chips, for a total of thirty channels of readout. Two ACDC boards are used for each LAPPD, to handle the full 56 channels of the photosensor, with 2 spare channels on each board. Signals can be coupled to the spare channels through SMA connectors on the pickup board.
@@
The ACDC separates the multiplexed beam gate and heartbeat signals on the basis of their distinct signal widths. Data from any LAPPD self-trigger that occurs within the configurable beam gate width (currently set to 20 s) is recorded. The ACDCs record timestamps for all received beam gate signals and all self-triggers. Short synchronization (sync) data frames record the local time stamp of the hearbeat signal. For most of the data-taking period, one in ten heartbeat signals were recorded (i.e. one sync frame every ten seconds).
@@ 4.1 Event Building
For all LAPPD data taken prior to September 2023, we apply an additional correction to partially compensate for instabilities in the phase-locked loop (PLL) synchronization with the central clock, that have since been mitigated. For data taken after that point, the effect was corrected in hardware. After synchronization offset fitting and the secondary correction, the time difference between an LAPPD data frame and the central beam trigger must be less than s for the LAPPD data frame to be matched. This is loose enough to account for both the programmed signal delay applied by the ACC and residual uncertainties in the LAPPD-CTC timing synchronization.
@@
Figure 8 shows, for the full sample of beam events, the difference between the time at which the LAPPD received the beam gate and the time at which the LAPPD self-triggered. The structure of this distribution shows a central peak around 1.6 s wide, corresponding to the beam spill, with a substantial ( events ns) and relatively flat background to either side. That background is a combination of dark and electronics noise triggers, cosmics, and beam-related backgrounds arising during the spill. The presence of a peak consistent with the beam spill indicates that the LAPPD is in fact detecting events correlated with the beam.

### v2

# First Light from Beam Neutrinos on an LAPPD in ANNIE

@@ Abstract
The Accelerator Neutrino Neutron Interaction Experiment (ANNIE) is both a physics experiment and a technology testbed for next-generation light-based neutrino detection. In this paper, we report the first demonstration of a fully integrated Large Area Picosecond Photodetector (LAPPD) operating in a running neutrino beam experiment. Particular focus is given to the design, commissioning, and successful deployment of the Packaged ANNIE LAPPD (PAL), a waterproof, self-triggering module incorporating fast waveform digitization and precision timing synchronized to the ANNIE detector subsystems. We identify beam-correlated LAPPD data frames consistent with charged-current neutrino interactions observed in multiple detector subsystems, establishing the first detection of neutrino-induced Cherenkov light with an LAPPD. These results validate the system-level performance of LAPPDs under realistic experimental conditions—including long-term stability, timing synchronization, and event matching with conventional PMT and muon detector systems—marking a critical step toward their deployment in future large-scale neutrino and particle detectors.
[…]
@@ 1 Introduction
LAPPDs are MicroChannel Plate (MCP)-based [mcpref] imaging photosensors with a variety of potential applications in high-energy and nuclear physics, as well as practical applications in medical imaging [LAPPDhist] and other areas. Produced using advanced fabrication techniques such as atomic layer deposition, LAPPDs offer gain [IncomPerformance], approximately 50 ps time resolution [timingpaper], and unprecedented large area coverage. At 20 cm x 20 cm, they have a form factor nine times larger than the next-largest MCP-PMTs, at significantly lower cost per unit area. The large LAPPD area is coupled with a unique ability to correlate photon arrival times with sub-cm position resolution, determined primarily by the structure of the anode.
Until this point, studies of LAPPDs have been limited to proof-of-principle studies on the bench-top, focused primarily on characterizing LAPPD response in the single-photoelectron regime. Investigating LAPPD performance in the context of a running experiment, using actual neutrino interactions, represents a critical milestone toward establishing LAPPDs as viable next-generation photodetector for particle physics applications. Here we present the first neutrino interactions detected using an LAPPD in a working experiment and document the relevant technical details of the system developed to further the application readiness of this technology. In Section 2, we briefly describe the ANNIE detector and the experimental context. In Sec. 3, we describe the integration of an LAPPD into the ANNIE experiment. Finally, in Section 4, we establish the first detection of beam neutrinos with an LAPPD.
@@ 2 Review of the ANNIE Experiment
ANNIE is both a physics experiment studying neutrino-nucleus interactions in a water target and a unique US platform for developing next-generation neutrino detection technology [ANNIEProposal]. The primary physics goal of ANNIE is to study neutrino interactions on water, with specific focus on measuring the abundance of final state neutrons from neutrino-nucleus interactions, using gadolinium (Gd) enhanced water. Measurements of final-state neutron multiplicity will improve our understanding of neutrino-nucleus interactions. The complex many-body physics of neutrino-nucleus scattering is among the dominant systematic uncertainties in precision long-baseline oscillation measurements [oscillationsE, oscillationsE2].
ANNIE also supports a broad R&D program that serves the experiment’s core physics goals while developing and testing new technologies for the future. One of the central goals of ANNIE is the first demonstration of LAPPDs in a neutrino experiment, as described in this paper. Other milestones of the R&D program include deployment of the first Gd-loaded water target in a neutrino beam (2021) and the first water-based liquid Scintillator (WbLS) target in a neutrino beam [ANNIE:2023yny].
@@ 2.1 The ANNIE Detector
Figure 1 shows the ANNIE detector configuration. The main target consists of an upright cylindrical steel tank filled with 26 tons of Gd-loaded (0.1% by mass) ultra-pure deionized water. The photodetectors inside the tank are a mix of 8-, 10- and 11-inch conventional photomultiplier tubes (PMTs) that provide coverage in all directions. Eight regularly spaced access ports around the lid of the water tank each permit a column of up to three LAPPDs to be inserted into the tank. Data presented in this paper was acquired with a single LAPPD at a central position on wall of the tank downstream of the beam. In addition to the water volume, the full ANNIE detector system also includes a downstream muon detection system (refurbished from SciBooNE’s Muon Range Detector, or MRD [SciBooNE:2008bzb]) that provides the primary constraint on muon energy and direction, as well as a scintillator-based Front Muon Veto (FMV) upstream of the water target.
A muon produced by a neutrino interaction in the fiducial volume will propagate through the tank, producing Cherenkov light along its path, before entering into and passing through the layers of scintillator paddles and iron absorber in the MRD. The position and timing of scintillator hits, combined with information about the Cherenkov light distribution from the tank photodetectors, are used to reconstruct the muon track. Neutrons produced by the neutrino interaction scatter and lose energy (thermalize) allowing them to be captured on either H or Gd in the active volume. The Gd loading substantially increases the neutron detection efficiency by increasing the capture cross-section for thermal neutrons (thereby shortening the neutron capture time) and providing a higher-energy, more easily-detected 8 GeV de-excitation -ray cascade [FSneutrons, EGADS].
@@ 2.2 Booster Neutrino Beam
The ANNIE detector and data acquisition system are designed around the Booster Neutrino Beam (BNB) [MiniBooNE:2008hfu], which delivers neutrinos in discrete bursts timed to the accelerator Radio Frequency (RF) structure. The low duty cycle of the BNB allows the experiment to statistically suppress continuum backgrounds by confining the readout of the various detector sub-systems to a short time window around the arrival of neutrinos. A subset of off-beam data provides a sample of the continuum backgrounds, while laser, LED, and radioactive sources are used for calibration.
BNB neutrinos are produced from the decay of mesons made by colliding 8 GeV protons with a stationary beryllium target located at the MI-12 facility. Pions in a narrow range of momenta are selected using magnetic horns and steered to a decay pipe, where they decay to form a beam of 93 pure muon neutrinos (with 6 muon anti-neutrino and 1 electron (anti) neutrinos) with an energy spectrum peaked at a little below 600 MeV [MiniBooNE:2008hfu].
@@ 2.3 System Timing and DAQ Software
The system clock for the entire ANNIE experiment is provided by a CAEN V1495 general purpose VME board with a customizable FPGA, dubbed the ANNIE Central Trigger Card (CTC) [V1495]. The CTC is disciplined to a 25 MHz clock signal provided by a customized Spectrum Instruments TM-4 GPS clock with a double-ovenized crystal oscillator [TM4]. The CTC multiplies up the 25 MHz input clock to 125 MHz and passes that clock to the various subsystems. All subsystems match their timing to this clock via phase-locked loops (PLLs). Where needed, the CTC forwards a heartbeat, or pulse-per-second (PPS) signal, also provided by the GPS clock, to the various subsystems.
The CTC also manages, forwards, and timestamps a variety of beam-related and calibration triggers. Two of these signals are of particular importance. The CTC forms a beam trigger or beam gate as the logical and of a fiber-optic copy of the Booster Extraction Sync (BES) signal which precedes beam extraction by 320 s, and the Tevatron clock $1D signal, which indicates that the beam will be accelerated and extracted to the BNB and hence delivered to ANNIE. It then forwards time-delayed copies of this beam trigger to the readout electronics for the conventional PMT readout system and MRD, triggering both systems to read out all data related to a spill.
@@
For DAQ software, ANNIE uses the ToolDAQ framework, developed at Warwick University for the HyperK collaboration [benajmin_richards_2018_1482767]. Asynchronous modular code sequences (ToolChains) pull data from each of the subsystems. An overall system loop on the top-level DAQ computer queries and pulls data from each subsystem ToolChain when available. CTC data, which includes timestamps for the beam input and heartbeat signals, is also pulled as a data stream. These time stamps are used to tie together the responses from individual subsystems to build a complete record of the detector’s response to a beam spill, cosmic ray, or calibration source, hereafter referred to as an event.
@@ 3 Technical Integration of an LAPPD into ANNIE
The design of this system was shaped by several key design requirements. Fast sampling electronics and synchronization of the LAPPDs and other subsystems with a reliable central clock are necessary for ANNIE’s ultimate goal of demonstrating 100ps global time stamping of photon hits. Power dissipation from the readout electronics needed to be minimized to reduce heating effects at the LAPPD and in the interior of the waterproof housing. This ensured stable operating conditions (the resistance and performance of the MCPs can be temperature-sensitive) and also minimized temperature-related physical stresses on the LAPPD. Signals and power were routed between the surface and the tank in a way that maximized signal fidelity and power stability while minimizing the risk of water leaks.
@@ 3.1 Surface Electronics
VME-based Annie Central Cards (ACCs) provide communications and clock synchronization with the underwater readout system, and serve as conduits for transferring the digital outputs of the LAPPDs to the ANNIE DAQ computer using a Raspberry Pi. Each ACC card is capable of servicing up to 2 LAPPDs. The ACCs are controlled by Altera Arria V FPGAs [ArriaV]. The communication link connecting the ACC and ACDC boards relies on the Low Voltage Differential Signaling (LVDS) logic standard. Each link between an ACC and ACDC consists of 4 twisted pairs: two lines for transmitting and receiving data, one line for sending the beam gate, and a clock signal for ensuring clock-domain alignment. A USB-serial buffer chip operating at 48 Mbps transfers data between the ACC and the control computer. While the ACC cards are designed to sit in a VME crate and have a standard VME form-factor, they do not require a VME backplane. They are housed in a VME crate with no backplane and powered by an Excelsys power supply[Excelsys].
@@ 3.2 Packaged ANNIE LAPPD (PAL)
A PAL consists of a waterproof housing containing the LAPPD assembly with readout electronics (ACDC cards), the slow controls board (LVHV), and independent leak-sensing and temperature-monitoring devices, as well as flood-proof connectors that mate to waterproof cables. These elements are pictured in Figure 3.
@@
The first LAPPD deployed in ANNIE, LAPPD-40 (numbered reflecting its production order by the manufacturer, Incom Inc.), is a first-generation model LAPPD [IncomPerformance] instrumented with 28 silk-screened silver anode microstrips that are integral with the module. Photon hits are reconstructed based on the arrival of signal pulses at the two ends of each microstrip, as detailed further in Ref [multiphoton]. The nominal gain of LAPPD-40, as measured by Incom, is at 950V per MCP, and the average Quantum Efficiency (QE) is 17. This performance is on the low end of the LAPPDs received by ANNIE, due in part to LAPPD-40 being one of the earlier functioning models.
@@ 3.2.1 Structure of the PAL
The physical waterproof housing of the PAL consists of a 2-inch thick PVC sidewall sandwiched between a stainless steel backplate and a quarter-inch thick UVT acrylic window. The window and backplate are bolted onto the sidewall with chemical-resistant Viton fluoroelastomer 3/16-inch O-rings [Viton] providing the waterproof seal. All of the power and communications lines connect to the LVHV board, which is affixed to the inside top of the sidewall. The LAPPD assembly is mounted onto tabs extending from the sidewall. All HV cables, power and communications lines from the ACDC cards and trigger cards plug into the LVHV board.
@@ 3.2.2 Digitization and Triggering in the PAL
Sampling and digitization of LAPPD waveforms is handled by ACDC cards, custom boards designed at the University of Chicago, that attach to the APB through high-density connectors. The ACDCs are built around the 6-channel PSEC4 waveform sampling ASIC, which provides 10 GS/s sampling rates and 1 GHz analog bandwidth [PSEC]. Each ACDC card contains five PSEC4 chips, for a total of thirty channels of readout. Two ACDC boards are used for each LAPPD, to handle the full 56 channels of the photosensor, with 2 spare channels on each board. Signals can be coupled to the spare channels through SMA connectors on the pickup board.
@@
The LAPPD readout electronics were bench-tested to verify that channel-to-channel timing alignment was better than the 100 ps sampling interval. Both in bench tests and in situ, a loss of synchronization between PSEC4 chips would manifest as a simultaneous 8 ns shift across five readout channels; no such occurrences were observed. Calibration with synthetic signals was used to measure and correct path-length differences in the PCB pickup board traces. Small biases of order 10 ps may remain, arising from non-uniform PSEC4 sampling rates, and will be investigated in future work. These nonuniformities fall well below our current target of 100 ps single-photon time resolution.
@@
The ACDC separates the multiplexed beam gate and heartbeat signals on the basis of their distinct signal widths. Data from any LAPPD self-trigger that occurs within the configurable beam gate width (currently set to 20 s) is recorded. The ACDCs record timestamps for all received beam gate signals and all self-triggers. Short synchronization (sync) data frames record the local time stamp of the heartbeat signal. For most of the data-taking period, one in ten heartbeat signals were recorded (i.e. one sync frame every ten seconds).
@@ 4.1 Event Building
A timing instability in the phase-locked loop (PLL) synchronization of the PAL with the central clock was identified during 2023 and fully resolved with a hardware update in September of that year. All data collected since that update exhibit stable timing and require no additional treatment. The data presented in this paper, however, were acquired prior to the fix and therefore require an additional software-based correction to compensate for clock drift. After applying the synchronization offset fit and the secondary correction, any LAPPD data frame occurring within s of the central beam time can be reliably matched to the rest of the system. This is loose enough to account for both the programmed signal delay applied by the ACC and residual uncertainties in the LAPPD-CTC timing synchronization.
@@
Figure 8 shows, for the full sample of beam events, the difference between the time at which the LAPPD received the beam gate and the time at which the LAPPD self-triggered. The structure of this distribution shows a central peak around 1.6 s wide, corresponding to the beam spill, with a substantial ( events ns) and relatively flat background to either side. That background is a combination of dark and electronics noise triggers, cosmic ray muons, and beam-related backgrounds arising during the spill. The presence of a peak consistent with the beam spill indicates that the LAPPD is in fact detecting events correlated with the beam.